# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ A.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $||f_{\text{mean}}(X)-f_{\text{mean}}(Y)||>L\,d_{M}(X,Y)$, which is a contradiction. Thus, the mean function is not Lipschitz continuous with respect to the matching distance.
A.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the mean function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, consisting of $3$ and $2$ vectors of dimension $d$, respectively. Then, we have $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{3}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[2]}||\mathbf{v}_{i}-\mathbf{u}_{j}||=||\mathbf{v}_{3}-\mathbf{u}_{2}||=\sqrt{d}\epsilon$. We also have that:
$\Big{|}\Big{|}f_{\text{mean}}(X)-f_{\text{mean}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\frac{1}{3}\sum_{i=1}^{3}\mathbf{v}_{i}-\frac{1}{2}\sum_{j=1}^{2}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\frac{1}{3}\mathbf{v}_{1}+\frac{1}{3}\mathbf{v}_{2}+\frac{1}{3}\mathbf{v}_{3}-\frac{1}{2}\mathbf{u}_{1}-\frac{1}{2}\mathbf{u}_{2}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\frac{1}{3}\mathbf{v}_{3}\bigg{|}\bigg{|}$
@@ A.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
For any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $||f_{\text{mean}}(X)-f_{\text{mean}}(Y)||>L\,d_{H}(X,Y)$. We have thus reached a contradiction. Therefore, the mean function is not Lipschitz continuous with respect to the Hausdorff distance.
A.3.4 The sum function is not Lipschitz continuous with respect to EMD
Suppose that the sum function is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Then, let $m=\lfloor L+1\rfloor$. Let also $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}\!\!\}$ be two multisets, each consisting of $m$ vectors. We construct the two sets such that $\mathbf{v}_{1}=\mathbf{u}_{1},\mathbf{v}_{2}=\mathbf{u}_{2},\ldots,\mathbf{v}_{m-1}=\mathbf{u}_{m-1}$, and $\mathbf{v}_{1}+\ldots+\mathbf{v}_{m-1}=\mathbf{u}_{1}+\ldots+\mathbf{u}_{m-1}=0$. Let also $||\mathbf{v}_{m}-\mathbf{u}_{m}||=m$. We already showed in subsection A.3.1 that the distance between the mean vectors of two multisets of vectors is a lower bound on the EMD between them. Therefore, we have that:
$\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}d_{ij}$ $\leq\bigg{|}\bigg{|}\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{m}\sum_{j=1}^{n}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\frac{1}{m}||(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})||$
$=\frac{1}{m}||\mathbf{v}_{m}-\mathbf{u}_{m}||$
@@ A.3.4 The sum function is not Lipschitz continuous with respect to EMD
$[\mathbf{F}]_{ij}=\begin{cases}1&\text{if }i=j\\ 0&\text{if }i\neq j\\ \end{cases}$
Therefore, the EMD between $X$ and $Y$ is equal to $1/m$. Then, we have that:
$\Big{|}\Big{|}f_{\text{sum}}(X)-f_{\text{sum}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\sum_{i=1}^{m}\mathbf{v}_{i}-\sum_{j=1}^{m}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=||(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})||$
$=||\mathbf{v}_{m}-\mathbf{u}_{m}||$
$=m\cdot 1\frac{1}{m}$
$=m\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}||\mathbf{v}_{i}-\mathbf{u}_{j}||$
$>L\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}||\mathbf{v}_{i}-\mathbf{u}_{j}||$
@@
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ such that $||f_{\text{sum}}(X)-f_{\text{sum}}(Y)||>L\,d_{\text{EMD}}(X,Y)$. We have thus arrived at a contradiction, and the sum function is not Lipschitz continuous with respect to EMD.
A.3.5 The sum function is Lipschitz continuous with respect to the matching distance
@@
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||+||\mathbf{v}_{\pi(n+1)}||+\ldots+||\mathbf{v}_{\pi(m)}||\bigg{]}\$
@@
$\Big{|}\Big{|}f_{\text{sum}}(X)-f_{\text{sum}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\sum_{i=1}^{m}\mathbf{v}_{i}-\sum_{j=1}^{n}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=||(\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1})+\ldots+(\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n})+\mathbf{v}_{\pi^{*}(n+1)}+\ldots+\mathbf{v}_{\pi^{*}(m)}||$
$\leq||\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}||+\ldots+||\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n}||+||\mathbf{v}_{\pi^{*}(n+1)}||+\ldots+||\mathbf{v}_{\pi^{*}(m)}||$
$=\sum_{i=1}^{n}||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||+||\mathbf{v}_{\pi^{*}(n+1)}||+\ldots+||\mathbf{v}_{\pi^{*}(m)}||$
$=\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||+||\mathbf{v}_{\pi(n+1)}||+\ldots+||\mathbf{v}_{\pi(m)}||\bigg{]}$
@@
A.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff Distance
Suppose that the sum function is Lipschitz continuous with respect to the Hausdorff Distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we have $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{2}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[2]}\min_{j\in[1]}||\mathbf{v}_{i}-\mathbf{u}_{j}||=||\mathbf{v}_{2}-\mathbf{u}_{1}||=\sqrt{d}\epsilon$. We also have that:
$\Big{|}\Big{|}f_{\text{sum}}(X)-f_{\text{sum}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\sum_{i=1}^{2}\mathbf{v}_{i}-\mathbf{u}_{1}\bigg{|}\bigg{|}$
$=||\mathbf{v}_{1}+\mathbf{v}_{2}-\mathbf{u}_{1}||$
$=||\mathbf{v}_{2}||$
@@ A.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff Distance
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ such that $||f_{\text{sum}}(X)-f_{\text{sum}}(Y)||>L\,d_{H}(X,Y)$, which is a contradiction. Therefore, the sum function is also not Lipschitz continuous with respect to the Hausdorff distance.
A.3.7 The max function is not Lipschitz continuous with respect to EMD
Suppose that the max function is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Then, let $m=\lfloor L+1\rfloor$. Let also $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}\!\!\}$ be two multisets, each consisting of $m$ $d$-dimensional vectors. We construct the two sets such that $\mathbf{v}_{1}=\mathbf{u}_{1},\mathbf{v}_{2}=\mathbf{u}_{2},\ldots,\mathbf{v}_{m-1}=\mathbf{u}_{m-1}$, and $\mathbf{v}_{1}+\ldots+\mathbf{v}_{m-1}=\mathbf{u}_{1}+\ldots+\mathbf{u}_{m-1}=0$. Suppose that the elements of vectors $\mathbf{v}_{m}$, $\mathbf{u}_{m}$ are larger than those of all other vectors of $X$ and $Y$, respectively. Therefore, we have that $[\mathbf{v}_{m}]_{k}\leq[\mathbf{v}_{i}]_{k}$, $\forall i\in[m]$ and $k\in[d]$. We also have that $[\mathbf{u}_{m}]_{k}\leq[\mathbf{u}_{j}]_{k}$, $\forall j\in[m]$ and $k\in[d]$. Let also $||\mathbf{v}_{m}-\mathbf{u}_{m}||=1$. We already showed in subsection A.3.1 that the distance between the mean vectors of two multisets of vectors is a lower bound on the EMD between them. Therefore, we have that:
$\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}||\mathbf{v}_{i}-\mathbf{u}_{j}||$ $\leq\bigg{|}\bigg{|}\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{m}\sum_{j=1}^{n}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\frac{1}{m}||(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})||$
$=\frac{1}{m}||\mathbf{v}_{m}-\mathbf{u}_{m}||$

### v3

@@ B.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\|>L\,d_{M}(X,Y)$, which is a contradiction. Thus, the mean function is not Lipschitz continuous with respect to the matching distance.
B.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the mean function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, consisting of $3$ and $2$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{3}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[2]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. We also have that:
$\Big\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\Big\|$ $=\bigg\|\frac{1}{3}\sum_{i=1}^{3}\mathbf{v}_{i}-\frac{1}{2}\sum_{j=1}^{2}\mathbf{u}_{j}\bigg\|$
$=\bigg\|\frac{1}{3}\mathbf{v}_{1}+\frac{1}{3}\mathbf{v}_{2}+\frac{1}{3}\mathbf{v}_{3}-\frac{1}{2}\mathbf{u}_{1}-\frac{1}{2}\mathbf{u}_{2}\bigg\|$
$=\bigg\|\frac{1}{3}\mathbf{v}_{3}\bigg\|$
@@ B.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
For any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\|>L\,d_{H}(X,Y)$. We have thus reached a contradiction. Therefore, the mean function is not Lipschitz continuous with respect to the Hausdorff distance.
B.3.4 The sum function is not Lipschitz continuous with respect to EMD
Suppose that the sum function is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Then, let $m=\lfloor L+1\rfloor$. Let also $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}\!\!\}$ be two multisets, each consisting of $m$ vectors. We construct the two multisets such that $\mathbf{v}_{1}=\mathbf{u}_{1},\mathbf{v}_{2}=\mathbf{u}_{2},\ldots,\mathbf{v}_{m-1}=\mathbf{u}_{m-1}$, and such that $\mathbf{v}_{1}+\ldots+\mathbf{v}_{m-1}=\mathbf{u}_{1}+\ldots+\mathbf{u}_{m-1}=0$. Let also $\|\mathbf{v}_{m}-\mathbf{u}_{m}\|=m$. We already showed in subsection B.3.1 that the distance between the mean vectors of two multisets of vectors is a lower bound on the EMD between them. Therefore, we have that:
$\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$ $\geq\bigg\|\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{m}\sum_{j=1}^{m}\mathbf{u}_{j}\bigg\|$
$=\frac{1}{m}\|(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})\|$
$=\frac{1}{m}\|\mathbf{v}_{m}-\mathbf{u}_{m}\|$
@@ B.3.4 The sum function is not Lipschitz continuous with respect to EMD
[…]
Therefore, the EMD between $X$ and $Y$ is equal to $1$. Then, we have that:
$\Big\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\Big\|$ $=\bigg\|\sum_{i=1}^{m}\mathbf{v}_{i}-\sum_{j=1}^{m}\mathbf{u}_{j}\bigg\|$
$=\|(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})\|$
$=\|\mathbf{v}_{m}-\mathbf{u}_{m}\|$
$=m\cdot 1$
$=m\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}^{*}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
$>L\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}^{*}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
@@
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\|>L\,d_{\text{EMD}}(X,Y)$. We have thus arrived at a contradiction, and the sum function is not Lipschitz continuous with respect to EMD.
B.3.5 The sum function is Lipschitz continuous with respect to the matching distance
@@
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|+\|\mathbf{v}_{\pi(n+1)}\|+\ldots+\|\mathbf{v}_{\pi(m)}\|\bigg]\\$
@@
$\Big\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\Big\|$ $=\bigg\|\sum_{i=1}^{m}\mathbf{v}_{i}-\sum_{j=1}^{n}\mathbf{u}_{j}\bigg\|$
$=\|(\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1})+\ldots+(\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n})+\mathbf{v}_{\pi^{*}(n+1)}+\ldots+\mathbf{v}_{\pi^{*}(m)}\|$
$\leq\|\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}\|+\ldots+\|\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n}\|+\|\mathbf{v}_{\pi^{*}(n+1)}\|+\ldots+\|\mathbf{v}_{\pi^{*}(m)}\|$
$=\sum_{i=1}^{n}\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|+\|\mathbf{v}_{\pi^{*}(n+1)}\|+\ldots+\|\mathbf{v}_{\pi^{*}(m)}\|$
$=\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|+\|\mathbf{v}_{\pi(n+1)}\|+\ldots+\|\mathbf{v}_{\pi(m)}\|\bigg]$
@@
B.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the sum function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{2}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[2]}\min_{j\in[1]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{2}-\mathbf{u}_{1}\|=\sqrt{d}\epsilon$. We also have that:
$\Big\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\Big\|$ $=\bigg\|\sum_{i=1}^{2}\mathbf{v}_{i}-\mathbf{u}_{1}\bigg\|$
$=\|\mathbf{v}_{1}+\mathbf{v}_{2}-\mathbf{u}_{1}\|$
$=\|\mathbf{v}_{2}\|$
@@ B.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\|>L\,d_{H}(X,Y)$, which is a contradiction. Therefore, the sum function is also not Lipschitz continuous with respect to the Hausdorff distance.
B.3.7 The max function is not Lipschitz continuous with respect to EMD
Suppose that the max function is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Then, let $m=\lfloor L+1\rfloor$. Let also $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}\!\!\}$ be two multisets, each consisting of $m$ $d$-dimensional vectors. We construct the two sets such that $\mathbf{v}_{1}=\mathbf{u}_{1},\mathbf{v}_{2}=\mathbf{u}_{2},\ldots,\mathbf{v}_{m-1}=\mathbf{u}_{m-1}$, and such that $\mathbf{v}_{1}+\ldots+\mathbf{v}_{m-1}=\mathbf{u}_{1}+\ldots+\mathbf{u}_{m-1}=0$. Suppose that the elements of vectors $\mathbf{v}_{m}$, $\mathbf{u}_{m}$ are larger than those of all other vectors of $X$ and $Y$, respectively. Therefore, we have that $\mathbf{v}_{m}]_{k}\geq[\mathbf{v}_{i}]_{k}$, $\forall i\in[m]$ and $k\in[d]$. We also have that $\mathbf{u}_{m}]_{k}\geq[\mathbf{u}_{j}]_{k}$, $\forall j\in[m]$ and $k\in[d]$. Let also $\|\mathbf{v}_{m}-\mathbf{u}_{m}\|=1$. We already showed in subsection B.3.1 that the distance between the mean vectors of two multisets of vectors is a lower bound on the EMD between them. Therefore, we have that:
$\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$ $\geq\bigg\|\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{m}\sum_{j=1}^{m}\mathbf{u}_{j}\bigg\|$
$=\frac{1}{m}\|(\mathbf{v}_{1}-\mathbf{u}_{1})+(\mathbf{v}_{2}-\mathbf{u}_{2})+\ldots+(\mathbf{v}_{m}-\mathbf{u}_{m})\|$
$=\frac{1}{m}\|\mathbf{v}_{m}-\mathbf{u}_{m}\|$